# Parallel Split Learning with Global Sampling

### v1

@@ V-C Time Complexity of LDS
We measured the total number of EM iterations invoked during LDS for various values of $\Delta$, $K$, $N$, and $R$. We present our results for $p_{s}=0.1$ and $p_{s}=0.2$, respectively, as a 3D heatmap in Figure 2. We observe that reinitialization ($R=1$) significantly impacts the number of iterations. Without reinitialization ($R=0$), the EM algorithm requires on average less than a third of the iterations to converge. This indicates that the $\boldsymbol{\pi}$ from the previous iteration provides a good starting point for the next iteration. The highest number of iterations occurs when using $R=1$ with a high value of $\Delta$. The EM algorithm needs more iterations to converge when the $\boldsymbol{\alpha}$ are adjusted more aggressively. Poor initial values due to reinitialization strongly exacerbates this effect. For a higher number of clients $K$, the number of iterations increased in all cases. This is primarily because the EM algorithm must be re-run each time a client dataset depletes. The results show that the choice of sample size $N$ has no significant impact on the number of iterations. This is due to the speedup achieved by the reformulation of the E-step in Section IV-C.
Lastly, we measure the total sampling time required for LDS and compare it with UGS. We use a straggler probability of $p_{s}=0.2$, a batch size of $B=128$, and different values of $\Delta$. The results, shown in Figure 3, indicate that the sampling time for LDS increases as $K$ increases. This is because the number of EM invokations and the number of computations for the $\Gamma_{k}$ scales with $K$. However, LDS is still only slightly slower than UGS, even for high values of $\Delta$. Thus, LDS not only provides a significant reduction in TPE but also maintains a low computational overhead compared to UGS.
@@ V-D Limitations
Our methods have several limitations. First, we assume that the delay times do not change significantly during an epoch. However, this assumption may not hold in practice. The straggler mitigation may not be effective if the delay times change frequently. Furthermore, the inherent reliance on gradient averaging causes increased latency and communication overhead. Gradient averaging exacerbates the straggler effect, as the server must wait for all contributing clients to send their gradients before proceeding. Lastly, we observed that when the number of clients is at least as large as the global batch size, the local batch sizes can become very small, which leads to increased latency due to idle clients. The integration of UGS and LDS with other advanced techniques like local loss function design and client clustering could further enhance the performance of PSL, by removing the need for gradient averaging and improving the scalability of the framework.
@@ VI Conclusion
In this paper, we proposed the UGS and LDS methods as an extension to the PSL framework introduced in [22]. Our methods addressed key challenges in PSL, including large effective batch sizes, Non-IID data distributions, and the straggler effect. UGS mitigates the large effective batch size problem by allowing variable local batch sizes, decoupling the number of clients from the effective batch size. Our analysis demonstrated that UGS effectively reduces batch deviations, even under Non-IID data settings. LDS is a generalization of UGS and balances the trade-off between batch deviation and training time using a single hyperparameter $\Delta$ that controls the trade-off. Our extensive simulations on the CIFAR10 dataset showed that PSL with UGS and LDS outperforms existing frameworks in terms of test accuracy and convergence stability. Specifically, our methods achieved a test accuracy close to the central learning baseline in the Non-IID setting, while other frameworks experienced significant drops in performance. Compared to PSL with fixed local batch sizes, UGS and LDS increased the test accuracy by up to $34.1\%$ in the Non-IID setting. The evaluation of LDS for straggler mitigation revealed substantial reduction in Time-Per-Epoch, achieving a reduction of up to $62\%$. Our findings suggest that UGS and LDS are promising solutions for enhancing PSL performance in real-world applications. Venues for future research include improving the client scalability, reducing the communication overhead, and decreasing the training latency of PSL.
@@ Appendix B Proofs
$=\text{Var}(Y^{\prime}_{m})+\left(\mathbb{E}[Y^{\prime}_{m}]-\mathbb{E}[Y_{m}]\right)^{2}$
@@
$\left|\frac{Y^{\prime}_{m}}{B}-\beta_{0,m}\right|\geq\epsilon\implies\left(\frac{Y^{\prime}_{m}}{B}-\beta_{0,m}\right)^{2}\geq\epsilon^{2}$
@@
$P\left(\left|\frac{Y^{\prime}_{m}}{B}-\beta_{0,m}\right|\geq\epsilon\right)\leq\frac{\text{Var}(Y^{\prime}_{m})+\left(\mathbb{E}[Y^{\prime}_{m}]-\mathbb{E}[Y_{m}]\right)^{2}}{B^{2}\epsilon^{2}}$
@@
Case 1: $\beta_{k,m}=\beta_{0,m}$.
@@
$=\beta_{0,m}B=\mathbb{E}[Y_{m}]$
Case 2: $B_{k}=B\frac{D_{k}}{D}$.
@@
$\iff$ $\mathbb{E}[Y^{\prime}_{m}]=\mathbb{E}[Y_{m}]$
@@
$P\left(\left|\frac{Y^{\prime}_{m}}{B}-\beta_{0,m}\right|\geq\epsilon\right)\leq\frac{\textnormal{Var}(Y^{\prime}_{m})}{B^{2}\epsilon^{2}}$
@@
Where Equation 6 follows from Jensen’s inequality with the concave function $\varphi(x)=x(1-x)$. Equality holds if and only if ${\beta_{k,m}=\beta_{0,m}}$ for all $k\in\mathcal{K}$. Thus, it holds that
@@
Where $B_{k}=B\pi_{k}=B\frac{D_{k}}{D}$. According to Theorem Theorem 1 the bound holds.
@@
We introduce a Lagrange multiplier $\lambda$ to ensure the constraint $\sum_{k\in\mathcal{K}}\pi_{k}=1$, which yields the following objective function:
$\mathcal{L}(\boldsymbol{\pi},\lambda)=\mathcal{Q}(\pi,\pi^{\text{old}})+\ln p(\pi\mid\boldsymbol{\alpha})+\lambda\left(\sum_{k\in\mathcal{K}}\pi_{k}-1\right)$
@@
$\iff$ $\pi_{k}=\frac{N_{k}+\alpha_{k}-1}{N+\alpha_{0}-K}$

### v2

@@ V-C Time Complexity of LDS
We measured the total number of EM iterations invoked during LDS for various values of $\Delta$, $K$, $N$, and $R$. We present our results for $p_{s}=0.1$ and $p_{s}=0.2$, respectively, as a 3D heatmap in Figure 2. We observe that reinitialization ($R=1$) significantly impacts the number of iterations. Without reinitialization ($R=0$), the EM algorithm requires on average less than a third of the iterations to converge. This indicates that the $\boldsymbol{\pi}$ from the previous iteration provides a good starting point for the next iteration. The highest number of iterations occurs when using $R=1$ with a high value of $\Delta$ and a small sample size. The EM algorithm needs more iterations to converge when the $\boldsymbol{\alpha}$ are adjusted more aggressively. In particular, with a small sample size. Poor initial values due to reinitialization strongly exacerbates this effect. For a higher number of clients $K$, the number of iterations increased in all cases. This is primarily because the EM algorithm must be re-run each time a client dataset depletes. The results show that for $R=0$ the choice of sample size $N$ has no significant impact on the number of iterations. This is due to the speedup achieved by the reformulation of the E-step in Section IV-C.
Lastly, we measured the total sampling time required for LDS and compared it with UGS. We used a straggler probability of $p_{s}=0.2$, a batch size of $B=128$, and different values of $\Delta$. The results, shown in Figure 3, indicate that the sampling time for LDS increases as $K$ increases. This is because the number of EM invokations and the number of computations for the $\hat{\boldsymbol{\gamma}}_{k}$ scales with $K$. However, LDS is still only slightly slower than UGS, even for high values of $\Delta$. Thus, LDS not only provides a significant reduction in TPE but also maintains a low computational overhead compared to UGS.
@@ V-D Limitations
We outline some limitations of our methods that were observed during our analysis and simulations. First, we assume that the delay times do not change significantly during an epoch. However, this assumption may not hold in practice. The straggler mitigation may not be effective if the delay times change frequently. Furthermore, the inherent reliance on gradient averaging causes increased latency and communication overhead. Gradient averaging exacerbates the straggler effect, as the server must wait for all contributing clients to send their gradients before proceeding. Lastly, we observed that when the number of clients is at least as large as the global batch size, the local batch sizes can become very small, which leads to increased latency due to idle clients. The integration of UGS and LDS with other advanced techniques like local loss function design and client clustering could further enhance the performance of PSL, by removing the need for gradient averaging and improving the scalability of the framework. Other venues for future research include reducing the communication overhead and decreasing the training latency of PSL.
@@ VI Conclusion
In this paper, we proposed the UGS and LDS methods as an extension to the PSL framework. Our methods addressed key challenges in PSL, including large effective batch sizes, non-IID data distributions, and the straggler effect. UGS mitigates the large effective batch size problem by allowing variable local batch sizes, decoupling the number of clients from the effective batch size. Our analysis demonstrated that UGS effectively reduces batch deviations, even under non-IID data settings. LDS is a generalization of UGS and balances the trade-off between batch deviation and training time using a single hyperparameter $\Delta$ that controls the trade-off. Our extensive simulations on the CIFAR10 dataset showed that PSL with UGS and LDS outperforms existing frameworks in terms of test accuracy and convergence stability. Specifically, our methods achieved a test accuracy close to the central learning baseline in the non-IID setting, while other frameworks experienced significant drops in performance. Compared to PSL with fixed local batch sizes, UGS and LDS increased the test accuracy by up to $34.1\%$ in the non-IID setting. The evaluation of LDS for straggler mitigation revealed substantial reduction in training time per epoch, achieving a reduction of up to $62\%$. Our findings suggest that UGS and LDS are promising solutions for enhancing PSL performance in real-world applications.
VII Acknowledgment
This work was supported by the Federal Ministry of Education and Research (BMBF, Germany) as part of NeuroSys: Impulse durch Anwendungen (Projekt D) - A under Grant 03ZU1106DA.
@@ Appendix B Proofs
$=\text{Var}(Y^{\prime}_{m})+\left(\mathbb{E}[Y^{\prime}_{m}]-\mathbb{E}[Y_{m}]\right)^{2},$
@@
$\left|\frac{Y^{\prime}_{m}}{B}-\beta_{0,m}\right|\geq\epsilon\implies\left(\frac{Y^{\prime}_{m}}{B}-\beta_{0,m}\right)^{2}\geq\epsilon^{2}.$
@@
$P\left(\left|\frac{Y^{\prime}_{m}}{B}-\beta_{0,m}\right|\geq\epsilon\right)\leq\frac{\text{Var}(Y^{\prime}_{m})+\left(\mathbb{E}[Y^{\prime}_{m}]-\mathbb{E}[Y_{m}]\right)^{2}}{B^{2}\epsilon^{2}}.$
@@
Case 1: $\beta_{k,m}=\beta_{0,m}:$
@@
$=\beta_{0,m}B=\mathbb{E}[Y_{m}].$
Case 2: $B_{k}=B\frac{D_{k}}{D}:$
@@
$\iff$ $\mathbb{E}[Y^{\prime}_{m}]=\mathbb{E}[Y_{m}].$
@@
$P\left(\left|\frac{Y^{\prime}_{m}}{B}-\beta_{0,m}\right|\geq\epsilon\right)\leq\frac{\textnormal{Var}(Y^{\prime}_{m})}{B^{2}\epsilon^{2}}.$
@@
where Inequation (6) follows from Jensen’s inequality with the concave function $\varphi(x)=x(1-x)$. Equality holds if and only if ${\beta_{k,m}=\beta_{0,m}}$ for all $k\in\mathcal{K}$. Thus, it holds that
@@
Where $B_{k}=B\pi_{k}=B\frac{D_{k}}{D}$. According to Theorem 1 the bound holds.
@@
We introduce a Lagrange multiplier $\lambda$ to ensure constraint $\sum_{k\in\mathcal{K}}\pi_{k}=1$, which yields the following objective function:
$\mathcal{L}(\boldsymbol{\pi},\lambda)=\mathcal{Q}(\boldsymbol{\pi},\boldsymbol{\pi}^{\text{old}})+\ln P(\boldsymbol{\pi}\mid\boldsymbol{\alpha})+\lambda\left(\sum_{k\in\mathcal{K}}\pi_{k}-1\right).$
@@
$\iff$ $\pi_{k}=\frac{N_{k}+\alpha_{k}-1}{N+\alpha_{0}-K}.$